OpenStudy (moongazer):

if f(x)= (x+1)/(x^2-1) and g(x) = (3x+7)/2x find f[g(x)]. Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

step on write \[f(g(x))\]

OpenStudy (anonymous):

step two replace \[g(x)\] by \[\frac{3x+7}{2x}\] to get \[f(\frac{3x+7}{2x})\]

OpenStudy (anonymous):

so far so good?

OpenStudy (moongazer):

yup

OpenStudy (anonymous):

then were we see an x in \[f(x)\] we replace it by \[\frac{3x+7}{2x}\] to get \[\frac{\frac{3x+7}{2x}+1}{(\frac{3x+7}{2x})^2-1}\]

OpenStudy (anonymous):

zat ok?

OpenStudy (anonymous):

is jsut an exercise in simplication

OpenStudy (anonymous):

simplification

OpenStudy (anonymous):

well actually why don't we simplify to begin with. \[f(x)=\frac{x+1}{x^2-1}=\frac{x+1}{(x+1)(x-1)}=\frac{1}{x-1}\]\]

OpenStudy (moongazer):

yup that is what my teacher told me

OpenStudy (anonymous):

so we could make this problem much easier if we use the second f i wrote. on the other hand this is not really valid

OpenStudy (anonymous):

those function are not identical. they are identical if x is not -1. if x is -1 the first function is undefined, whereas the second function is \[\frac{1}{-2}\]

OpenStudy (anonymous):

but if your teacher does not mind, neither do i. just write \[f(x)=\frac{1}{x-1}\] and so \[f(g(x))=f(\frac{3x+7}{2x})=\frac{1}{\frac{3x+7}{2x}-1}\]

OpenStudy (moongazer):

what??

OpenStudy (anonymous):

then multiply top and bottom by 2x to get \[f(g(x))=\frac{2x}{3x+7-2x}=\frac{2x}{x+7}\]\][

OpenStudy (moongazer):

what happened to the ( )^2?

OpenStudy (anonymous):

oops i lost you

OpenStudy (anonymous):

are we going to replace \[f(x)=\frac{x+1}{x^2-1}\] by \[f(x)=\frac{1}{x-1}\]?

OpenStudy (moongazer):

yes because i think it is easier

OpenStudy (anonymous):

ok in that case there is no square in the denominator

OpenStudy (moongazer):

ooooohhh now i know

OpenStudy (anonymous):

if we make that change in f, then \[f(g(x))=f(\frac{3x+7}{2x})\] yes?

OpenStudy (anonymous):

if so then everything i wrote above works yes?

OpenStudy (anonymous):

if not let me know

OpenStudy (moongazer):

isn't it it should first be 1/(3x+7)- 1?

OpenStudy (moongazer):

am i correct?

OpenStudy (anonymous):

no.

OpenStudy (moongazer):

why?

OpenStudy (anonymous):

\[g(x)=\frac{3x+7}{2x}\]

OpenStudy (anonymous):

not \[3x+7\]

OpenStudy (anonymous):

so when we replace x in f(x) by \[\frac{3x+7}{2x}\] we get \[\frac{1}{\frac{3x+7}{2x}-1}\]

OpenStudy (moongazer):

yup i forgot to type 2x sorry for that

OpenStudy (moongazer):

please continue

OpenStudy (anonymous):

ok good

OpenStudy (anonymous):

now you can simplify this compound fraction by multiplying top and bottom by 2x

OpenStudy (anonymous):

we have \[f(g(x))=\frac{1}{\frac{3x+7}{2x}-1}\] \[=\frac{1}{\frac{3x+7}{2x}-1}\times \frac{2x}{2x}\] \[=\frac{2x}{3x+7-2x}\] \[=\frac{2x}{x+7}\]

OpenStudy (anonymous):

k?

OpenStudy (moongazer):

what happened with the 2x below 3x+7?

OpenStudy (anonymous):

i multiplied top and bottom by 2x to get rid of it. canceled from the denominator

OpenStudy (anonymous):

\[\frac{3x+7}{2x}\times 2x=3x+7\]

OpenStudy (moongazer):

ohhh ok

OpenStudy (anonymous):

you cannot have a compound fraction as an answer, so i got rid of it by multiplying top and bottom by that annoying 2x in the denominator of the denominator

OpenStudy (anonymous):

clear or no?

OpenStudy (moongazer):

wait ill do it on a paper ^_^

OpenStudy (anonymous):

k

OpenStudy (moongazer):

how did it become 2x/2x+7?

OpenStudy (anonymous):

ok let us start with \[\frac{1}{\frac{3x+7}{2x}-1}\] ok?

OpenStudy (moongazer):

ok

OpenStudy (anonymous):

multiply top and bottom by 2x. what do you get?

OpenStudy (anonymous):

\[1\times 2x=2x\] that is your numerator

OpenStudy (moongazer):

ok

OpenStudy (anonymous):

\[\frac{3x+7}{2x}\times 2x = 3x+7\] that is why you did it to get rid of the pesky 2x in the bottom

OpenStudy (anonymous):

\[-1\times 2x=-2x\]

OpenStudy (moongazer):

ok

OpenStudy (anonymous):

of course you have to multiply the whole denominator by 2x, not just part of it.

OpenStudy (anonymous):

so you numerator is 2x

OpenStudy (anonymous):

your denominator is 3x+7-2x

OpenStudy (anonymous):

which give a denominator of x+7

OpenStudy (moongazer):

ok

OpenStudy (anonymous):

so our "final answer" is \[f(g(x))=\frac{2x}{x+7}\]

OpenStudy (moongazer):

i just got confused i forgot that 3x and -2x can be combined

OpenStudy (anonymous):

ah but clear now yes?

OpenStudy (moongazer):

yes! thanks^_^

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

ah but clear now yes?